Homework #4 Cost of Capital 2. Vo = So + Bo, where V. is the firm value in period 0, S, is the stockholders equity in periodHomework #4 Cost of Capital Debt $15 Debt $17 The Firm The Firm Time 0 to Time 1 Assets = $100 Assets = $110 Rate of Return?
Answer #1

A. Rate of return for the firm = (V1-V0)/V0

=(110-100)/100 =10%

B. Rate of Return for the debt = (B1-B0)/B0

=(17-15)/15 =13.33%

C. S0 = V0- B0 = 100-15 =85

S1= V1- B1 = 110-17 = 93

D. Rate of return for equity = (S1-S0)/S0

= (93-85)/85 =9.412%

E. Return for the firm =10%

Return for the firm on weighted average basis

= [(1/3)×13.33]+[(2/3)×9.412] =10% (Rounded off)